# Forecasting and Manipulating the Forecasts of Others

### 18 Mar 2026 v2

@@ D.2 Deviation propagation
A spike perturbation of trader $i$’s demand at time $t$ enters aggregate order flow, triggering a coupled forward response. For $s>t$, let $\delta D_{\Sigma,s}$ be the aggregate demand perturbation and $\delta w_{s}^{j}(u)$ the noise-state density perturbation for observer $j$. The deviation system (cf. Section B.2) is:
$\delta D_{\Sigma,s}$ $=\sum_{k\neq i}\int_{0}^{s}D^{k}_{s}(u)\,\delta w_{s}^{k}(u)\,du,$ (D.2)
$\partial_{s}\delta w_{s}^{0}(u)$ $=\widetilde{C}^{0}(s,u)^{\top}\biggl(\sigma_{0}^{-1}\delta D_{\Sigma,s}-\int_{0}^{s}C^{0}(s,r)\,\delta w_{s}^{0}(r)\,dr\biggr),$ (D.3)
@@
$\partial_{s}\delta w_{s}^{k}(u)$ $=\widetilde{C}^{k}(s,u)^{\top}\biggl(\begin{pmatrix}\sigma_{0}^{-1}\delta D_{\Sigma,s}\\ 0\end{pmatrix}-\int_{0}^{s}C^{k}(s,r)\,\delta w_{s}^{k}(r)\,dr\biggr),\quad k\neq i,$ (D.4)
$\delta P_{s}$ $=\sigma_{V}\int_{0}^{s}E^{V}\delta w_{s}^{0}(u)\,du,$ (D.5)
@@
with $\delta w_{t}^{j}\equiv 0$, $\delta D_{\Sigma,t}=1$, $\delta I_{s}^{i}=0$. The demand perturbation enters each observer’s stacked observation through the order-flow row only. Equation (D.2) closes the loop: opponents’ noise-state shifts feed through their policy kernels back into aggregate demand. Since $E^{V}E^{0,\top}=0$, $\delta P_{t}=\lambda(t)$; when $n=1$, $\delta D_{\Sigma,s}=0$ for $s>t$ and $\delta P_{s}=\lambda(t)$ for all $s\geq t$.
D.3 First-order condition
The optimality condition from a unit demand impulse at $t$ is
$\mathbb{E}\!\left[V_{T}-P_{t}-\int_{t}^{T}D_{\tau}^{i}\,\delta P_{\tau}\,d\tau\;\Big|\;\mathcal{F}_{t}^{i}\right]=0.$ (D.6)
The mispricing is $\mathbb{E}[V_{T}-P_{t}\mid\mathcal{F}_{t}^{i}]=\sigma_{V}\int_{0}^{t}E^{V}\,d_{u}(\widehat{W}_{t}^{i}-\widehat{W}_{t}^{0})(u)$ (since $\mathcal{F}_{t}^{0}\subset\mathcal{F}_{t}^{i}$). The conditional future demand in noise-state linear form is $\mathbb{E}[D_{\tau}^{i}\mid\mathcal{F}_{t}^{i}]=\int_{0}^{t}D^{i}_{\tau}(u)\,d_{u}\widehat{W}_{t}^{i}(u)$ for $\tau>t$ (the noise-state is a martingale, so only $u\leq t$ contributes). Both terms are linear in $\widehat{W}_{t}^{i}$, so matching noise-state densities gives (5.3). $D^{i}_{t}(u)$ does not appear directly: the integral over $[t,T]$ does not isolate $\tau=t$. To pin down $D^{i}_{t}(u)$, we differentiate both sides in $t$.
D.4 Leibniz differentiation and backward ODE
Impact adjoint.
Leibniz on $H^{i,\mathrm{imp}}$ gives
$\frac{d}{dt}H^{i,\mathrm{imp}}_{t}(u)=-\lambda(t)\,D^{i}_{t}(u)+\int_{t}^{T}\partial_{t}\delta P_{\tau}\,D^{i}_{\tau}(u)\,d\tau,$ (D.7)
where the boundary term $\delta P_{t}\cdot D^{i}_{t}(u)=-\lambda(t)D^{i}_{t}(u)$ is instantaneous self-impact—this is where $\lambda(t)$ enters as the endogenous $G^{DD,i}$ and where $D^{i}_{t}(u)$ first appears explicitly.
Value adjoint.
In primitive-noise coordinates (Theorem 3.1(i)), $H^{i,\mathrm{val}}_{t}(u)$ has a $t$-independent term (the direct observation gap $\Pi_{\mathcal{E}}^{i}-\Pi^{0}$) and an integral $\int_{0}^{t}(F_{t}^{i}-F_{t}^{0})(v,u)\,dv$. Differentiating by Leibniz and the kernel evolution $\partial_{t}F_{t}^{j}(v,u)=\widetilde{C}^{j}(t,v)^{\top}\widetilde{C}^{j}(t,u)$:
@@ D.4 Leibniz differentiation and backward ODE
$\frac{d}{dt}H^{i,\mathrm{val}}_{t}(u)=\sigma_{V}E^{V}\biggl[\bigl(F_{t}^{i}-F_{t}^{0}\bigr)(t,u)+\int_{0}^{t}\bigl(\widetilde{C}^{i}(t,v)^{\top}\widetilde{C}^{i}(t,u)-\widetilde{C}^{0}(t,v)^{\top}\widetilde{C}^{0}(t,u)\bigr)\,dv\biggr].$ (D.8)
Closed backward ODE.
Equating $\frac{d}{dt}H^{i,\mathrm{val}}=\frac{d}{dt}H^{i,\mathrm{imp}}$ and solving for $D^{i}_{t}(u)$ yields (5.4). The derivative $\partial_{t}\delta P_{\tau}$ is obtained by differentiating (D.2)–(D.5) in the impulse time, following the same logic as Section B.2 with $\sqrt{P^{k}}X\to C^{k}$, $\widetilde{X}^{k}\to\widetilde{C}^{k}$. Convexity (from $\lambda(t)>0$) ensures sufficiency.

### 23 Mar 2026 v3

@@ E.1 The price-impact kernel and Kyle’s lambda
A unit spike perturbation of trader $i$’s demand at time $t$ triggers a coupled forward response: the market maker revises the price, each opponent revises their estimate, adjusts their trading, and the adjustments feed back into prices. Let $\delta D_{\Sigma,\tau}$ denote the aggregate demand perturbation and $\delta w_{\tau}^{j}(u)$ the noise-state density perturbation for observer $j$. The deviation system is:
$\delta D_{\Sigma,\tau}$ $=\sum_{k\neq i}\int_{0}^{\tau}D^{k}_{\tau}(u)\,\delta w_{\tau}^{k}(u)\,du,$ (E.1)
$\partial_{\tau}\delta w_{\tau}^{0}(u)$ $=\widetilde{C}^{0}(\tau,u)^{\top}\biggl(\sigma_{0}^{-1}\delta D_{\Sigma,\tau}-\int_{0}^{\tau}C^{0}(\tau,r)\,\delta w_{\tau}^{0}(r)\,dr\biggr),$ (E.2)
@@
$\partial_{\tau}\delta w_{\tau}^{k}(u)$ $=\widetilde{C}^{k}(\tau,u)^{\top}\biggl(\begin{pmatrix}\sigma_{0}^{-1}\delta D_{\Sigma,\tau}\\ 0\end{pmatrix}-\int_{0}^{\tau}C^{k}(\tau,r)\,\delta w_{\tau}^{k}(r)\,dr\biggr),\quad k\neq i,$ (E.3)
@@
with initial conditions $\delta w_{t}^{j}\equiv 0$, $\delta D_{\Sigma,t}=1$, and $\delta I_{\tau}^{i}=0$ (the deviator’s own innovation is unaffected). Equation (E.1) closes the loop: each opponent’s noise-state shift propagates through their policy kernel back into aggregate demand, which in turn shifts the market maker’s and other opponents’ noise-states.
Price-impact kernel.
Define the price-impact kernel as the price response to a unit demand impulse at time $t$:
$l(\tau,t):=\sigma_{V}\int_{0}^{\tau}E^{V}\delta w_{\tau}^{0}(u)\,du,\qquad\tau\geq t.$ (E.4)
This is the object $\ell(t,s)$ of the price decomposition (5.1): the market maker’s filtering gains are deterministic, so $l(\tau,t)$ is deterministic.
Kyle’s lambda.
Since $E^{V}E^{0,\top}=0$, the market maker learns about the fundamental only through drift-based inference, and the instantaneous price impact is
$l(t,t)=\lambda(t):=\sigma_{V}\!\int_{0}^{t}E^{V}\widetilde{C}^{0}(t,u)^{\top}du.$ (E.5)
E.2 Leibniz differentiation and the backward ODE
Differentiating (5.3) in $t$, Leibniz on the impact side produces $\lambda(t)\,D^{i}_{t}$ plus $\mathbb{E}[\int_{t}^{T}\frac{\partial l}{\partial t}(\tau,t)\,D^{i}_{\tau}\,d\tau\mid\mathcal{F}_{t}^{i}]$. The value side requires more care.
Value side.
The conditional mispricing $H^{i,\mathrm{val}}_{t}=\widehat{V}_{t}^{i}-P_{t}=\sigma_{V}E^{V}\bigl(\widehat{W}^{i}_{t}(t)-\widehat{W}^{0}_{t}(t)\bigr)$ can be written as
$\sigma_{V}E^{V}\int_{0}^{t}\biggl[\int_{0}^{t}\bigl(F^{i}_{t}(u,s)-F^{0}_{t}(u,s)\bigr)\,du\biggr]dW_{s}\,.$
The integrand kernel is deterministic and evolves through $\partial_{t}F_{t}^{j}(v,u)=\widetilde{C}^{j}(t,v)^{\top}\widetilde{C}^{j}(t,u)$. Differentiating by Leibniz:
$dH^{i,\mathrm{val}}_{t}=\sigma_{V}E^{V}\int_{0}^{t}\biggl[\int_{0}^{t}\bigl(\widetilde{C}^{i}(t,u)^{\top}\widetilde{C}^{i}(t,s)-\widetilde{C}^{0}(t,u)^{\top}\widetilde{C}^{0}(t,s)\bigr)\,du\biggr]dW_{s}\,dt\;+\;dM_{t}\,$, (E.6)
where $dM_{t}$ collects the martingale increment from the boundary at $s=t$. The remaining boundary term (from the upper limit of the inner integral) vanishes because $E^{V}E^{i,\top}=E^{V}E^{0,\top}=0$: neither player’s observation noise correlates with the fundamental.
To condition on $\mathcal{F}_{t}^{i}$, note that $\widetilde{C}^{i}(t,\cdot)$ is by definition the component of the drift kernel that player $i$ has not resolved—it is orthogonal to $\mathcal{F}_{t}^{i}$. The $\widetilde{C}^{i}$ cross terms therefore vanish under conditioning, and $dW_{s}$ projects to $d\widehat{W}^{i}_{t}(s)$, giving
@@ E.2 Leibniz differentiation and the backward ODE
$\mathbb{E}\!\left[\frac{dH^{i,\mathrm{val}}_{t}}{dt}\;\Big|\;\mathcal{F}_{t}^{i}\right]=-\sigma_{V}E^{V}\int_{0}^{t}\biggl[\int_{0}^{t}\widetilde{C}^{0}(t,u)^{\top}\widetilde{C}^{0}(t,s)\,du\biggr]d\widehat{W}^{i}_{t}(s)\,.$